Find all the 5th roots of (5-1)
1
Expert's answer
2021-01-04T19:07:04-0500

Let

a=5-i

IaI="\sqrt{ {5}^2+({-1})^2}\\"

IaI="\sqrt{26}\\"

"\alpha = tan^{-1}(-5)\\
a=|a|e^{i\alpha}=|a|(cos\alpha+isin\alpha)\\
a^{\frac{1}{5}}=|a|^{\frac{1}{5}}e^{i\frac{\alpha}{5}}\\
a^{\frac{1}{5}}=\sqrt{26}^{\frac{1}{5}}(cos\frac{tan^{-1}(\frac{1}{-5})+2kπ}{5}+isin\frac{tan^{-1}(\frac{1}{-5})+2kπ}{5})\\

k= 0, 1, 2, 3, 4"